FML Company in the U.S. has just signed a contract to purchase light rail cars from a manufacturer in Germany for euro 2,500,000. The purchase was made in June with payment in euro due six months later in December. Assume that in December, the axchange rate turns out to be as the company anticipated. If so, then the money market hedge turns out to be ____ than the unhedged position by ______.

The spot exchange rate is $1.40/euro

The six month forward rate is $1.38/euro

FML's weighted avg cost of capital is 11%

The Euro zone borrowing rate is 9%

The Euro zone lending rate is 7%

The US borrowing rate is 8%

The US lending rate is 6%

December call option and put option for euro 625,000; strike price $1.42, premium 1.5%

FML forecast for the spot rate six months from now is $1.43/euro.
Answer #1

Payment to be made in 6 months = Euro 2,500,000.

Spot Exchange Rate = $ 1.4/Euro and the 6 month Forward rate = $ 1.38/Euro

Under a money market hedge, the company will borrow in USD an amount equal to the discounted value (present value of Euro 2.5 million discounted at Euro zone lending rate) and convert the amount to Euro and place this amount in an interest bearing deposit in Euro. This Euro deposit, in 6 months, will become Euro 2.5 million which will be used for repayment. The mechanics will be as below:

Present Value of Euro payable = 2,500,000 / (1+7%)6/12? = Euro 2,416,841.22

FML will borrow in USD an amount equal to the above Euro amount converted at spot rate. Hence it will borrow (2416841.22 * 1.40) USD 3,383,577.71. The borrowing rate in USD is 8%. After 6 months, the company would have to pay (3383577.71 * (1+8%)6/12?) = USD 3,516,317.1

Now FML, will use the borrowed USD amount, convert into Euro 2,416,841.22 and place it in a Euro deposit at 7%. This deposit in 6 months will become Euro 2.5 million and will be used to repay the German manufacturer.

Hence the exchange rate under money market hedge for FML turns out to be (3516317.1/2500000) = $1.4065/Euro. Thus the monery market hedge turns out to be cheaper (better off) compared to being unhedged by $ 58682.8952 [(2,500,000 * 1.43) - 3516317.1]